# SplashNet: Split-and-Share Encoders for Accurate and Efficient Typing with Surface Electromyography

### v1

@@ Abstract
Surface electromyography (sEMG) at the wrists could enable natural, keyboard-free text entry, yet the state-of-the-art emg2qwerty baseline still misrecognizes 51.8% of characters zero-shot on unseen users and 7.0% after user-specific fine-tuning. We trace much of these errors to mismatched cross-user signal statistics, fragile reliance on high-order feature dependencies, and the absence of architectural inductive biases aligned with the bilateral nature of typing. To address these issues, we introduce three simple modifications: (i) Rolling Time Normalization which adaptively aligns input distributions across users; (ii) Aggressive Channel Masking, which encourages reliance on low-order feature combinations more likely to generalize across users; and (iii) a Split-and-Share encoder that processes each hand independently with weight-shared streams to reflect the bilateral symmetry of the neuromuscular system. Combined with a five-fold reduction in spectral resolution (33$\rightarrow$6 frequency bands), these components yield a compact Split-and-Share model, SplashNet-mini, which uses only ¼ the parameters and 0.6× the FLOPs of the baseline while reducing character error rate (CER) to 36.4% zero-shot and 5.9% after fine-tuning. An upscaled variant, SplashNet (½ parameters, 1.15× FLOPs of the baseline), further lowers error to 35.7% and 5.5%, representing 31% and 21% relative improvements in the zero-shot and finetuned settings, respectively. SplashNet therefore establishes a new state-of-the-art without requiring additional data.
@@ 1 Introduction
Although promising, EMG interfaces face critical challenges in generalization. EMG interfaces suffer from substantial domain shift across users and sessions: anatomy, electrode placement, fatigue, and day-to-day physiology all alter the signal for the same action (14). The emg2qwerty benchmark probes this shift with (i) a zero-shot test, where one trains on many users and tests on a new one, and (ii) a personalization test that finetunes on a few target-user samples. Their baseline ASR-style system (CNN on spectrograms + LM decoding) misrecognized $51.8\%$ of characters zero-shot, but fine-tuning cut errors to approximately $7\%$, showing that most residual errors are systematic, user-specific variations. Yet even large training pools struggle: Sivakumar et al. (14) estimate $\mathcal{O}(10^{3})$ users are needed for robust out-of-the-box performance, a trend echoed by CTRL-Labs’ gains from massive data (2).
@@ 2 Insights into improving sEMG decoding
We summarily make three insights, and subsequently, three simple modifications to sEMG decoding that substantially improve zero-shot and fine-tuning performance while reducing computational costs.
@@
Insight 2: sEMG feature subsets may be conserved across users. Through visualizing sEMG activity, we empirically find that while two typists rarely share identical high-dimensional EMG signatures for the same keystroke, they often share subsets of electrodes or frequency bands that behave similarly. A representative example is shown in Figure 1b, where we present the PSTHs of sEMG activity from three users. While the full set of features looks fairly different across users, randomly selected subsets of features have a greater chance of being similar between some users (Augmented View 1 and 2 in Figure 1b). Decoding strategies that hinge on these low-order combinations therefore stand a better chance of transferring across users. While there are many ways to encourage this, we focus on substantially modifying the hyperparameters of dataset augmentation to achieve this, incurring no test-time computational expense. In particular, we do Agressive Channel Masking (ACM) that on average zeros out more than half of the electrodes and removes broad spectral chunks from the rest. ACM encourages the model to rely on smaller, more universal feature sets (Figure 1b), which empirically improves zero-shot generalization.
Insight 3: an inductive bias for bilateral typing. Keyboard typing is fundamentally bilateral: each keystroke is driven solely by the muscles of its own wrist, while the neuromuscular mapping from activation to finger motion is almost mirror-symmetric across hands. We therefore propose a Split-and-Share architecture that accounts for these two facts in the model. Separate, but weight-shared, subnets encode left- and right-hand sEMG streams in parallel, converging only at the final linear layer where a single vocabulary-level softmax decodes the character (Figure 2). This enforces hand-specific locality, but prevents the encoder from entangling spurious cross-hand correlations. Further, re-use of parameters reduces resource demands: the compact Split-and-Share model uses 0.25× the parameters and approximately $60\%$ of the FLOPs of a conventional joint-hand encoder while matching zero-shot accuracy.
@@ 3 Related Works
Per-Session Normalization in EMG: The EMG and brain-computer interface (BCI) literature commonly normalize input signals prior to decoding to account for variability. For instance, prior EMG studies expressed amplitudes as a percentage of a reference contraction (maximum voluntary contraction) or applied z-score normalization based on a calibration recording (15). These normalizations enable comparisons across muscles and subjects (3; 8), and several works have normalized EMG features per session for machine learning pipelines (15; 10; 7; 20; 16; 6). In BCI literature, Willett et al. (19) show that continuously updating z-score statistics on spiking features within each session is essential for their high-performance speech neuroprosthesis, and Jarosiewicz et al. (5) used online bias correction to improve a point-and-click BCI. As described in Section 2, an important insight is not whether sEMG activity should be normalized, but how it is normalized. We ultimately find RTN significantly outperforms batch normalization.
@@
Multi-stream architectures and weight sharing: The concept of processing multiple information streams independently before later integration has been established in Automatic Speech Recognition (ASR), where multi-stream systems effectively handle speech at various resolutions or from multiple arrays (9). Similarly, in Sign Language Recognition (SLR), previous works have implemented separate processing streams for each hand (11; 17). Leveraging anatomical symmetry through weight sharing has shown success in pose estimation (21) and SLR (17). In the domain of EMG-based hand-pose recognition, Salter et al. (13) combined data from both hands into a unified dataset for single-model training, though this addresses an inherently unimanual task. For bimanual keystroke decoding, past methods have processed signals from both hands jointly (14; 18; 1). Our Split-and-Share architecture keeps the two wrists separate to respect biomechanics, yet ties the weights, effectively training a single-hand encoder with data from both hands. To our knowledge, this has not been explored in EMG keyboard decoding.
@@ 4.1 Reduced Spectral Granularity
Let $x_{c}(t)$ denote the raw 2 kHz EMG signal from electrode $c$. Following Sivakumar et al. (14) we first form a log-power spectrogram
@@
using a 64-point FFT ($33$ linearly spaced bins up to 1 kHz) and a hop size of 16. We instead aggregate the 33 bins into six broader, roughly log-spaced bands, using the same frequency ranges as in (2). We refer to this aggregation as reduced spectral granularity (RSG).
@@ 4.2 Rolling Time Normalization
Sivakumar et al. (14) normalizes each electrode’s spectrogram with batch-norm over the entire mini-batch, frequency, and time axes. This collapses all features recorded on an electrode into a single distribution and, crucially, re-uses statistics gathered during training at inference time—an issue when the test session (or user) drifts from the training distribution. RTN replaces this with a causal, per-feature normalizer that is computed independently for every sample, band, electrode channel and frequency bin. Let
@@
denote the log-power value at time-step $t$, mini-batch index $n$, band $b$, electrode $c$ and frequency bin $f$. RTN maintains cumulative statistics
@@
which are cheap to update via running sums in streaming settings. During a warm-up period of $T_{w}=125$ frames (the first $1$ second) the statistics are frozen to those computed over the entire warm-up window. The normalized output is
@@
where $\tilde{\mu}_{t}$ and $\tilde{\sigma}_{t}$ denote the warm-up-frozen statistics for $t<T_{w}$ and the cumulative statistics above for $t\geq T_{w}$. RTN adapts continuously to session-specific non-stationarities, avoids batch-level or training-level statistics, and is causal.
@@ 4.4 EMG Encoder Architectures
The TDSConv stack (4) is composed of alternating temporal convolution and fully connected blocks. Each convolutional block first reshapes the input from $\mathbb{R}^{T\times N\times D}$ to $\mathbb{R}^{N\times K\times H\times T}$, where $D=K\cdot H$ denotes the total feature dimension, $K$ is the number of convolution channels, and $H$ is the per-channel hidden width. A 2-D convolution with kernel size $1\times w$ is then applied along the time axis:
@@
Because the kernel height is 1, the same weights are reused at every hidden-width position $h$. The convolution still mixes all $K$ input channels to produce each of the $K$ output channels, giving the layer $K^{2}w$ parameters. The output is passed through ReLU, summed with the residual, and normalized using LayerNorm. The subsequent fully connected block consists of two linear layers with a ReLU in between and a residual connection:
@@
In the architecture of Sivakumar et al. (14), which we refer to as the Joint-Hand architecture, both hands are processed jointly following the Rotation-Invariant MLP, yielding a post-MLP embedding of $7682D=$ 768. The TDSConv stack operates with this full dimensionality throughout. In contrast, both the Split-only and Split-and-Share models operate on embeddings of $D=384$ per hand, and apply the TDSConv blocks separately to each hand’s input. This reduces the width of all fully connected layers from $768\times 768$ to $384\times 384$, leading to a four-fold reduction in parameters per FC layer, while preserving expressiveness by keeping the number of convolution channels $K=24$ and kernel width $w=32$ unchanged. Despite duplicating the encoder, the Split-only model has lower total FLOPs—approximately 60% of the baseline—due to the narrower per-hand embeddings. The Split-and-Share model uses the same dual-stream structure but shares all encoder weights between hands. This model has roughly half the parameters of the Split-only model, since almost all parameters (excluding the final prediction layer) are shared between the two encoders. Notably, the Split-only and Split-and-Share models have the same FLOPs.
@@ 4.5 Dataset Splits
The official emg2qwerty protocol introduced by Sivakumar et al. (14) assigns 100 participants to the training pool and eight to an unseen test pool. For the 96 training participants with $\geq 4$ recording sessions, every session except the final two is used to train generic models, while those last two sessions form the authors’ validation set; models are selected on this set before being evaluated on the eight held-out participants.
Because the validation data come from the same individuals seen during training, the procedure encourages models that memorize participant-specific idiosyncrasies rather than ones that generalize. The baseline model of Sivakumar et al. (14) illustrates this failure mode: with greedy decoding, it reaches a character-error rate (CER) of only $22.51\%$ on the training participants yet $55.38\%$ on the test participants. This indicates substantial overfitting.
To obtain a validation signal that better predicts zero-shot performance, we instead exploit the four training-pool participants who recorded $<4$ sessions and were therefore excluded from model fitting in the original setup. We take the final session of each of these four participants and combine them into what we term the other domain validation set. These participants exhibit notably noisier EMG data than those in the official test pool, providing a harder proxy for real-world generalization. Nonetheless, they provide a way of validating zero-shot generalization without requiring a complete restructuring of the official dataset splits.
@@ 5.1 Zero-shot model performance
Table 1: Zero-shot CER (%, mean $\pm$ s.d. across participants), GFLOPs, and parameters.
Method Other domain val Test domain val Test domain test GFLOPs (30 s) Params
Sivakumar et al. 2024 $52.10\pm 5.54$ $51.78\pm 4.61$ 201.325 5.29M
+ RSG $67.48$ $47.26\pm 5.26$ $47.18\pm 5.19$ 178.948 4.96M
+ RTN $61.95$ $39.49\pm 7.45$ $39.15\pm 6.20$ 178.948 4.96M
+ ACM $63.08$ $42.62\pm 7.18$ $42.62\pm 7.10$ 178.948 4.96M
+ RTN + ACM $58.85$ $36.41\pm 7.21$ $36.42\pm 7.11$ 178.948 4.96M
+ Split $58.64$ $37.28\pm 6.91$ $37.37\pm 7.34$ 118.346 2.68M
      + Share (SplashNet-mini) $58.20$ $36.46\pm 7.09$ $36.41\pm 7.30$ 118.346 1.38M
       + Upscale (SplashNet) $\bf{56.95}$ $\bf{35.49\pm 7.56}$ $\bf{35.67\pm 6.79}$ 229.384 2.58M
We report the performance of our proposed simple augmentations on the emg2qwerty zero-shot benchmark. Our analysis centers on beam-search decoding using the same 6-gram character LM as in Sivakumar et al. (14); corresponding greedy-decoding results appear in Appendix A.3. All reported p-values were obtained from one-tailed paired t-tests across participants. First, we apply RSG to replace the 33-bin spectrogram with six coarser frequency bands. RSG leads to a modest but consistent improvement in CER from $51.78\%$ to $47.18\%$ ($p=7e$-$5$). Part of this performance improvement may stem from the implicit amplification of frequency masking under SpecAugment (due to the lower spectral resolution).
Second, we replace batch-level normalization with RTN. RTN significantly improves zero-shot generalization ($p=6e$-$4$), reducing the CER to $39.15\%$. This suggests that input scale and shift differences across users are a primary obstacle to cross-user generalization in EMG decoding.
[…]

### v2

# SplashNet: Split‑and‑Share Encoders for Accurate and Efficient Typing with Surface Electromyography

@@ Abstract
Surface electromyography (sEMG) at the wrists could enable natural, keyboard‑free text entry, yet the state‑of‑the‑art emg2qwerty baseline still misrecognizes 51.8% of characters zero‑shot on unseen users and 7.0% after user‑specific fine‑tuning. We trace much of these errors to mismatched cross‑user signal statistics, fragile reliance on high‑order feature dependencies, and the absence of architectural inductive biases aligned with the bilateral nature of typing. To address these issues, we introduce three simple modifications: (i) Rolling Time Normalization which adaptively aligns input distributions across users; (ii) Aggressive Channel Masking, which encourages reliance on low‑order feature combinations more likely to generalize across users; and (iii) a Split‑and‑Share encoder that processes each hand independently with weight‑shared streams to reflect the bilateral symmetry of the neuromuscular system. Combined with a five‑fold reduction in spectral resolution ($3\rightarrow$6 frequency bands), these components yield a compact Split-and-Share model, SplashNet‑mini, which uses only ¼ the parameters and 0.6× the FLOPs of the baseline while reducing character error rate (CER) to 36.4% zero‑shot and 5.9% after fine‑tuning. An upscaled variant, SplashNet (½ parameters, 1.15× FLOPs of the baseline), further lowers error to 35.7% and 5.5%, representing 31% and 21% relative improvements in the zero-shot and finetuned settings, respectively. SplashNet therefore establishes a new state-of-the-art without requiring additional data.
@@ 1 Introduction
Although promising, EMG interfaces face critical challenges in generalization. EMG interfaces suffer from substantial domain shift across users and sessions: anatomy, electrode placement, fatigue, and day-to-day physiology all alter the signal for the same action (14). The emg2qwerty benchmark probes this shift with (i) a zero-shot test, where one trains on many users and tests on a new one, and (ii) a personalization test that finetunes on a few target-user samples. Their baseline ASR-style system (CNN on spectrograms + LM decoding) misrecognized $51.8\%$ of characters zero-shot, but finetuning cut errors to approximately $7\%$, showing that most residual errors are systematic, user-specific variations. Yet even large training pools struggle: Sivakumar et al. (14) estimate $\mathcal{O}(10^{3})$ users are needed for robust out-of-the-box performance, a trend echoed by CTRL-Labs’ gains from massive data (2).
@@ 2 Insights into improving sEMG decoding
We summarily make three insights, and subsequently, three simple modifications to sEMG decoding that substantially improve zero-shot and finetuning performance while reducing computational costs.
@@
Insight 2: sEMG feature subsets may be conserved across users. Through visualizing sEMG activity, we empirically find that while two typists rarely share identical high‑dimensional EMG signatures for the same keystroke, they often share subsets of electrodes or frequency bands that behave similarly. A representative example is shown in Figure 1b, where we present the PSTHs of sEMG activity from three users. While the full set of features looks fairly different across users, randomly selected subsets of features have a greater chance of being similar between some users (Augmented View 1 and 2 in Figure 1b). Decoding strategies that hinge on these low‑order combinations therefore stand a better chance of transferring across users. While there are many ways to encourage this, we focus on substantially modifying the hyperparameters of dataset augmentation to achieve this, incurring no test-time computational expense. In particular, we do Aggressive Channel Masking (ACM) that on average zeros out more than half of the electrodes and removes broad spectral chunks from the rest. ACM encourages the model to rely on smaller, more universal feature sets (Figure 1b), which empirically improves zero-shot generalization. We emphasize that this ‘low-order similarity’ perspective is a working hypothesis rather than a definitive statement about the structure of EMG representations: the empirical improvements we observe are consistent with this interpretation but do not prove it.
Insight 3: an inductive bias for bilateral typing. Keyboard typing is fundamentally bilateral: each keystroke is driven solely by the muscles of its own wrist, while the neuromuscular mapping from activation to finger motion is almost mirror‑symmetric across hands. We therefore propose a Split‑and‑Share architecture that accounts for these two facts in the model. Separate, but weight‑shared, subnets encode left‑ and right‑hand sEMG streams in parallel, converging only at the final linear layer where a single vocabulary‑level softmax decodes the character (Figure 2). This enforces hand‑specific locality, but prevents the encoder from entangling spurious cross‑hand correlations. Further, re-use of parameters reduces resource demands: the compact Split‑and‑Share model uses 0.25× the parameters and approximately $60\%$ of the FLOPs of a conventional joint‑hand encoder while matching zero‑shot accuracy.
@@ 3 Related Works
Per‑Session Normalization in EMG: The EMG and brain-computer interface (BCI) literature commonly normalize input signals prior to decoding to account for variability. For instance, prior EMG studies expressed amplitudes as a percentage of a reference contraction (maximum voluntary contraction) or applied z-score normalization based on a calibration recording (15). These normalizations enable comparisons across muscles and subjects (3; 8), and several works have normalized EMG features per session for machine learning pipelines (15; 10; 7; 20; 16; 6). In BCI literature, Willett et al. (19) show that continuously updating z‑score statistics on spiking features within each session is essential for their high‑performance speech neuroprosthesis, and Jarosiewicz et al. (5) used online bias correction to improve a point-and-click BCI. As described in Section 2, an important insight is not whether sEMG activity should be normalized, but how it is normalized. We ultimately find RTN significantly outperforms batch normalization.
@@
Multi-stream architectures and weight sharing: The concept of processing multiple information streams independently before later integration has been established in Automatic Speech Recognition (ASR), where multi-stream systems effectively handle speech at various resolutions or from multiple arrays (9). Similarly, in Sign Language Recognition (SLR), previous works have implemented separate processing streams for each hand (11; 17). Leveraging anatomical symmetry through weight sharing has shown success in pose estimation (21) and SLR (17). In the domain of EMG-based hand-pose recognition, Salter et al. (13) combined data from both hands into a unified dataset for single-model training, though this addresses an inherently unimanual task. For bimanual keystroke decoding, past methods have processed signals from both hands jointly (14; 18; 1). Our Split‑and‑Share architecture keeps the two wrists separate to respect biomechanics, yet ties the weights, effectively training a single-hand encoder with data from both hands. To our knowledge, this has not been explored in EMG keyboard decoding.
@@ 4.1 Reduced Spectral Granularity
Let $x_{c}(t)$ denote the raw 2 kHz EMG signal from electrode $c$. Following Sivakumar et al. (14) we first form a log‐power spectrogram
@@
using a 64‐point FFT ($33$ linearly spaced bins up to 1 kHz) and a hop size of 16. We instead aggregate the 33 bins into six broader, roughly log‐spaced bands, using the same frequency ranges as in (2). We refer to this aggregation as reduced spectral granularity (RSG).
@@ 4.2 Rolling Time Normalization
Sivakumar et al. (14) normalizes each electrode’s spectrogram with batch‑norm over the entire mini‑batch, frequency, and time axes. This collapses all features recorded on an electrode into a single distribution and, crucially, re‑uses statistics gathered during training at inference time—an issue when the test session (or user) drifts from the training distribution. RTN replaces this with a causal, per‑feature normalizer that is computed independently for every sample, band, electrode channel and frequency bin. Let
@@
denote the log‑power value at time‑step $t$, mini‑batch index $n$, band $b$, electrode $c$ and frequency bin $f$. RTN maintains cumulative statistics
@@
which are cheap to update via running sums in streaming settings. During a warm‑up period of $T_{w}=125$ frames (the first $1$ second) the statistics are frozen to those computed over the entire warm‑up window. The normalized output is
@@
where $\tilde{\mu}_{t}$ and $\tilde{\sigma}_{t}$ denote the warm‑up‑frozen statistics for $t<T_{w}$ and the cumulative statistics above for $t\geq T_{w}$. RTN adapts continuously to session-specific non-stationarities, avoids batch-level or training-level statistics, and is causal.
@@ 4.4 EMG Encoder Architectures
The TDSConv stack (4) is composed of alternating temporal convolution and fully connected blocks. Each convolutional block first reshapes the input from $\mathbb{R}^{T\times N\times D}$ to $\mathbb{R}^{N\times K\times H\times T}$, where $D=K\cdot H$ denotes the total feature dimension, $K$ is the number of convolution channels, and $H$ is the per‑channel hidden width. A 2‑D convolution with kernel size $1\times w$ is then applied along the time axis:
@@
Because the kernel height is 1, the same weights are reused at every hidden‑width position $h$. The convolution still mixes all $K$ input channels to produce each of the $K$ output channels, giving the layer $K^{2}w$ parameters. The output is passed through ReLU, summed with the residual, and normalized using LayerNorm. The subsequent fully connected block consists of two linear layers with a ReLU in between and a residual connection:
@@
In the architecture of Sivakumar et al. (14), which we refer to as the Joint-Hand architecture, both hands are processed jointly following the Rotation-Invariant MLP, yielding a post-MLP embedding of $2D=768$. The TDSConv stack operates with this full dimensionality throughout. In contrast, both the Split-only and Split-and-Share models operate on embeddings of $D=384$ per hand, and apply the TDSConv blocks separately to each hand’s input. This reduces the width of all fully connected layers from $768\times 768$ to $384\times 384$, leading to a four-fold reduction in parameters per FC layer, while preserving expressiveness by keeping the number of convolution channels $K=24$ and kernel width $w=32$ unchanged. Despite duplicating the encoder, the Split-only model has lower total FLOPs—approximately 60% of the baseline—due to the narrower per-hand embeddings. The Split-and-Share model uses the same dual-stream structure but shares all encoder weights between hands. This model has roughly half the parameters of the Split-only model, since almost all parameters (excluding the final prediction layer) are shared between the two encoders. Notably, the Split-only and Split-and-Share models have the same FLOPs.
@@
4.5 Backspace-aware beam search with 6-gram Character LM
For decoding we adopt the backspace‑aware beam search of Sivakumar et al. (14), which keeps the 50 most‑probable CTC label prefixes at each frame. Blank transitions simply update the score of an unchanged prefix, while the highest‑scoring non‑blank characters extend each hypothesis and are re‑ranked with a 6-gram character LM prior. When a backspace symbol appears, the algorithm retracts the last language‑model contribution, letting deletions correct earlier mistakes without extra passes. After the final frame any open LM context is closed, and the best‑scoring prefix provides the decoded keystroke sequence. All beam-search parameters are the same as in Sivakumar et al. (14)
@@ 4.6 Dataset Splits
The official emg2qwerty protocol introduced by Sivakumar et al. (14) assigns 100 participants to the training pool and eight to a held-out test pool. For the 96 training participants with $\geq 4$ recording sessions, every session except the final two is used to train generic models, while those last two sessions form the authors’ validation set; models are selected on this set before being evaluated on the eight held-out participants.
Because the validation data come from the same individuals seen during training, this procedure encourages models that memorize participant-specific idiosyncrasies rather than ones that generalize across users. To illustrate this effect while limiting computational cost, we evaluate training domain validation performance using the validation sessions from 18 of the 96 training users, which we term the training domain validation set. The same 18 users are also used in Appendix A.8, where they are fully held out from training to provide a more appropriate validation signal.
To obtain a validation signal that better predicts zero-shot performance in the main experiments, we instead exploit the four training-pool participants who recorded <4<4 sessions and were therefore excluded from model fitting in the original setup. We take the final session of each of these four participants and combine them into what we term the other domain validation set. These participants exhibit noisier EMG data than those in the official test pool, providing a harder proxy for real-world generalization. Nonetheless, they provide a way of validating zero-shot generalization without requiring a complete restructuring of the official dataset splits.
@@ 5.1 Zero-shot model performance
Table 1: Zero-shot CER (%, mean $\pm$ s.d. across participants), GFLOPs, and parameters. Columns in gray correspond to training domain validation results, which are reported for transparency but not used as indicators of generalization.
Method Train domain val
Other
domain
val
Test
domain
val
Test
domain
test
GFLOPs
(30 s)
Params
Sivakumar et al. 2024 $12.14$ $72.07$ $52.10\pm 5.54$ $51.78\pm 4.61$ 61.61 5.29M
+ RSG $13.52$ $67.48$ $47.26\pm 5.26$ $47.18\pm 5.19$ 54.15 4.96M
+ RTN $13.09$ $61.95$ $39.49\pm 7.45$ $39.15\pm 6.20$ 54.15 4.96M
+ ACM $23.47$ $63.08$ $42.62\pm 7.18$ $42.62\pm 7.10$ 54.15 4.96M
+ RTN + ACM $21.71$ $58.85$ $36.41\pm 7.21$ $36.42\pm 7.11$ 54.15 4.96M
+ Split $23.93$ $58.64$ $28\pm 6.91 37.37$$37\pm 7.34$ 36.84 2.68M
   + Share (SplashNet-mini) $26.44$ $58.20$ $36.46\pm 7.09$ $36.41\pm 7.30$ 36.84 1.38M
   + Upscale (SplashNet) $20.59$ $\bf{56.95}$ $\bf{35.49\pm 7.56}$ $\bf{35.67\pm 6.79}$ 71.38 2.58M
Before turning to the main cross-user generalization results, we first highlight a key discrepancy between training domain validation and held-out user evaluation. As shown in Table 1, the CERs on the training domain validation set are not only lower, but the relative ranking of models follows a markedly different pattern than on held-out user evaluation sets. This illustrates why training domain validation provides a misleading signal for cross-user generalization and should not be used for model selection. In Appendix A.8, we further show that when these same 18 users are held out entirely, their CERs rise substantially and align closely with the held-out evaluation trends shown here.
Our analysis centers on beam-search decoding using the same 6-gram character LM as in Sivakumar et al. (14); corresponding greedy-decoding results appear in Appendix A.3. All reported p-values were obtained from one-tailed paired t-tests across participants.
First, we apply RSG to replace the 33-bin spectrogram with six coarser frequency bands. RSG leads to a modest but consistent improvement in CER from $51.78\%$ to $47.18\%$ (p=7ep=7e-55). Part of this performance improvement may stem from the implicit amplification of frequency masking under SpecAugment (due to the lower spectral resolution).
Second, we replace batch-level normalization with RTN. RTN significantly improves zero-shot generalization (p=6ep=6e-44), reducing the CER to $39.15\%$. This suggests that input scale and shift differences across users are a primary obstacle to cross-user generalization in EMG decoding.
Third, we apply ACM, which encourages the model to rely on lower-order combinations of input features. Without RTN, ACM reduces zero-shot CER to $42.6\%$ ($p=.02$ vs. +RSG only). Combined with RTN, ACM reduces zero-shot CER to $36.42\%$ (p=5ep=5e-33 vs. +RSG+RTN), providing a strong Joint-Hand baseline.
Fourth, we explore architectural modifications that better reflect the causal and bilateral structure of EMG typing. We evaluate a Split-only model that encodes each hand separately without parameter sharing, achieving a CER of $37.37\%$. Despite having only half the parameters and $66\%$ of the FLOPs of our Joint-Hand baseline, this model performs competitively, though its lack of shared parameters may limit data efficiency. We then evaluate SplashNet-mini, where both hands are encoded via identical weight-shared encoders. This model achieves a $36.41\%$ CER—on par with the Joint-Hand baseline—while using just a quarter of its parameters and $66\%$ of the FLOPs. Finally, by increasing the embedding width and expanding the final convolutional layers, we create SplashNet, with similar FLOPs to the baseline of Sivakumar et al. (14) but still half the parameters. This yields a further improvement, reducing the CER to $35.67\%$ ($p=.049$ vs. SplashNet-mini).